OpenStudy (anonymous):

Area between the curves: y=(x-2)^2 ; y=x

OpenStudy (anonymous):



OpenStudy (amistre64):

determine the intersection points if they are not defined the integrate the difference of the functions across the interval

OpenStudy (anonymous):

(1,1) & (4,4)

OpenStudy (amistre64):

\[\int_{a}^{b}(x-2)^2-x~dx=\left.\frac13(x-2)^3-\frac12x^2\right|_{a}^{b}\]

OpenStudy (anonymous):

So I need the area between 1 and 4?

OpenStudy (amistre64):

yes

OpenStudy (phi):

although I think amistre has the bottom and top curves flipped

OpenStudy (amistre64):

i might; i tend to ignore the sign

OpenStudy (amistre64):

5-2=3 2-5=-3 since the area is an absolute value; the sign is useless

OpenStudy (anonymous):

\[A = -\frac{ 1 }{ 3 }x ^{3}+ \frac{ 5 }{ 2}x ^{2} - 4x | 4, 1\]

OpenStudy (amistre64):

\[\frac13(4-2)^3-\frac124^2-(\frac13(1-2)^3-\frac121^2)\] \[\frac83-8-(-\frac13-\frac12)\] \[\frac83-8+\frac13+\frac12\] is what i get

OpenStudy (anonymous):

Mine is different than yours, I expanded the (x-2)^2 and than figured it all together. Was this wrong?

OpenStudy (amistre64):

if you expanded correctly, then your way was fine

OpenStudy (anonymous):

I must of did something wrong, your answer is correct except for the sign.

OpenStudy (amistre64):

lol, story of me life ;)

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

So I can either keep the equations the way they are and Intergrate each one OR I can combine them as long as I do it correctly?

OpenStudy (amistre64):

(x−2)^2−x x^2-4x-x+4 x^2-5x+4 \[\frac13x^3-\frac52x^2+4x\]

OpenStudy (amistre64):

yes, the "doing it correctly" part is key ;)

OpenStudy (anonymous):

Ain't that the truth. Also, story of my life. I am understanding the Calculus for the most part but the Algebra is killing me at times.

OpenStudy (amistre64):

\[\frac134^3-\frac524^2+4(4)-(\frac131^3-\frac521^2+4(1))\] \[\frac{64}3-\frac{40}2+16-(\frac13-\frac52+4)\] \[\frac{64}3-20+16-\frac13+\frac52-4\] \[\frac{63}3-8+\frac52\] with any luck thats right too

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (amistre64):

good luck :) and yeah, it the algebra that bites